OpenStudy (anonymous):

through (2,6) and perpendicular to y=-5/4 + 1 (find the equaiton for the line)

OpenStudy (lgbasallote):

where's the x in y = -5/4 + 1?

OpenStudy (anonymous):

after -5/4 y=-5/4x + 1